The refractive index of germanium is 4.05. Find the speed of light in the medium and hence identify the medium
1
Expert's answer
2019-01-14T08:54:11-0500

The speed of light in the medium is

v=c/n=300000000/4.05=74100000 m/s.

The medium is germanium.